the length of rectangle is four more than twice of its weiht.Find the length and weight if the perimeter is 60cm.
Answer by CubeyThePenguin(3113)

width = w
length = 2w + 4

perimeter = 2(w) + 2(2w + 4) = 60
2w + 4w + 8 = 60
6w = 52

width = w = 26/3
length = 2w + 4 = 64/3